John gets scores of 88 and 78 on his first two tests. What score must he make to get an average grade of B, which is 80 points or greater?
I got 74.
Could someone please work this problem out so I could check my answer.
I'm unsure if I did everything correctly.

Found 2 solutions by Theo, greenestamps:
Answer by Theo(13342)

i think you're right.
let x = the third test score.
you get (88 + 78 + x) / 3 >= 80
multiply both sides of this equaiton by 3 to get:
88 + 78 + x >= 240
solve for x to get:
x >= 74.

when x = 74, the average is (88 + 78 + 74) / 3 = 80.

when x is greater than 74, the average will be greater than 80.

Answer by greenestamps(13198)



Here is an alternative method for solving problems like this involving the average of numbers that are "close together".

His first test is 8 points over the desired average.
His second test is 2 points under the desired average.
So his point total for the first two tests is 8-2 = 6 points over the desired average.
So to get EXACTLY an 80 average his next test should be 6 points under the desired average.

ANSWER: 80-6 = 74